Status Report

SOHO Daily Meeting Minutes for Thursday, 27 June 2002

By SpaceRef Editor
June 27, 2002
Filed under , ,

* RHESSI target for today is AR 0008; target for tomorrow
is TBD.

* The next talk in the Dynamics of Flux Ropes Lecture
Series will be today at 11am in Building 26, room G-10.
The topic will be: “Lagrangian Description of a Flux Rope.”

* Visitors will come to the EOF today around 3pm and tomorrow
around 11am.

* SPWG Friday at 2pm in the EAF. Please get calendar changes
to the SOC as soon as possible.

* SSR dump on Friday will be delayed until the end of the D16
pass at about 22:15 UT.

* Commanding times for next week:

Monday – D27 all day.
Tuesday – Commanding is good all day.
Wedesday – Gap until 17 UT then commanding is good.
Thursday – Gap from 16:30-19 UT then commanding is good.
Friday – Command on D66 until 17 UT then on D27.
Saturday – Command on D27 all day.
Sunday – Gap until 15 UT then commanding is good.

Spacecraft Status: Nominal

Spacecraft Anomalies: None

Accomplished Activities: VIRGO, SWAN

Planned Activities: NOM HGA, VIRGO, CELIAS

Upcoming Operations: 7/1 ACU/CDMU Memory Dumps

Ground Anomalies:
177/1840 6.25 hours no LMC data. D16. SCP problem. No impact
to operations.

06/26 DOY 177 – 06/27 DOY 178

EIT daily On-Line Solar Status and observations:

LASCO daily On-Line Solar Status and observations:

Web Page for Planning:

CDS: Nominal. Polarity reversal study at the south pole and
RHESSI support. Tomorrow – Active region study and
possibly a coronal hole study in the north.

UVCS: Nominal. 24-hour synoptic. Tomorrow – Cassini support and

LASCO: Nominal. C2/C3 synoptics, JOP159 support.

EIT: Nominal. 195 CME watch and synoptics.

MDI: Nominal. Full disk magnetograms and dopplergrams.

TRACE: Nominal. RHESSI support observing in UV on AR0008.

SpaceRef staff editor.